The max magnetic field strength of an electromagnetic field is 5 x 10^-6 T. Calculate the maximum electric field strength if the wave is travelling in a medium in which the speed of the wave is 0.75c.

1
Expert's answer
2022-01-05T04:21:42-0500

Solution

Given data

max magnetic field strength of an electromagnetic field is

"B_{max}=5 \times 10^{-6} T"

Velocity

"v=0.75\times3 \times 10^{8} m/sec"

The maximum strength of electric field is given by

"E_{max}=B_{max}v"

Putting all values

"E_{max}=5 \times 10^{-6} \times0.75\times3 \times 10^{8}"

.so

"E_{max}=1125V\/m"